# Assessing and improving reliability of neighbor embedding methods: a map-continuity perspective

### v1

@@ A.2.1 Perturbation score
In general, perturbation scores are not sensitive to perturbation length. Supplementary Figure S4 illustrates the perturbation scores of the CIFAR10 deep learning feature data for three perturbation lengths ($\lambda\in\{1,2,3\}$). Points with high perturbation scores remain consistent across different perturbation lengths. In practice, we recommend that users run perturbation score on a subset of data points and test with a few different perturbation lengths. Conceptually, the perturbation score detects points that fall within a radius of $\lambda$ around the location of the OI discontinuity.
@@
Moreover, we provide two approximation algorithms to accelerate the calculation of perturbation score for t-SNE along with a strategy for users to pre-screen points for which the perturbation score should be computed.
@@
For high-dimensional input data, often PCA as a pre-processing step is implemented before calculating the similarity scores. As similarity scores are recalculated for each perturbation we consider, PCA is repeated numerous times, leading to a significant increase of computation. Since PCA is robust to perturbing a single input point, we reuse the pre-processed input points after one PCA calculation based on the original input data. This approximation avoids multiple calculation of PCA. We find that this approximation is sufficiently accurate, as the differences between perturbation scores by approximation method 1 and the exact perturbation scores are empirically negligible (Supplementary Figure S6a).
@@
Given the input data matrix ${\bm{X}}=[{\bm{x}}_{1},\ldots,{\bm{x}}_{n}]^{\top}$ and perplexity $\mathcal{P}$, the computation of (exact) similarity scores $(v_{i,j}({\bm{X}}))_{i<j}$ in the t-SNE algorithm follows the steps below.
@@
    Calculate the pairwise distance $d_{ij}=\|{\bm{x}}_{i}-{\bm{x}}_{j}\|_{2}$ for $i,j=1,\ldots,n$.
@@
    $v_{i,j}({\bm{X}})=\frac{p_{j|i}+p_{i|j}}{2n}.$
The main computational bottleneck is at step 2, where we conduct binary search algorithm for $n$ times to solve $(\sigma_{i})_{1\leq i\leq n}$.
To provide an approximation method, we note that when perturbing the $k$-th point, for $i\neq k$, Eqn. 7 still approximately holds for the original standard deviation $\sigma_{i}$ since only one of the terms has been changed. Therefore, we can set $\widetilde{\sigma}_{i}\approx\sigma_{i}$ for $i\neq k$ as an approximation to $(\widetilde{\sigma}_{i})_{1\leq i\leq n}$, the standard deviations after perturbation. In this way, we only need to conduct the binary search once to solve $\widetilde{\sigma}_{k}$, which significantly speeds up the calculation of the similarity scores after perturbation.
In terms of computational performance, approximation method 2 leads to a reduction of running time by nearly $80\%$ for a dataset of size $5000$. We also find that approximation method 2 is highly accurate. As shown in Supplementary Figure S6b, perturbation scores based on approximation method 2 are approximately equal to the exact perturbation scores for most of the points.
@@
To further speed up the computation, we use the heuristic that embedding points receiving high perturbation scores are often found at the peripheries of clusters. This heuristic motivates us to calculate the perturbation scores only for the peripheral points in the embedding space, since they are mostly likely the unreliable points. We find that applying this pre-screening step tends to find most of the unreliable points (Supplementary Figure S5) with significantly increased computational speed.
@@
We use function dbscan in the R package dbscan (version 1.2-0) to identify embeddings on the periphery of clusters.
A.2.2 Singularity score
@@
Given an input data matrix ${\bm{X}}=[{\bm{x}}_{1},\ldots,{\bm{x}}_{n}]^{\top}$ and its embedding matrix ${\bm{Y}}=[{\bm{y}}_{1},\ldots,{\bm{y}}_{n}]^{\top}$, we describe our derivation of singularity scores. If we add an infinitesimal perturbation $\epsilon{\bm{e}}$ to ${\bm{x}}_{i}$, then by the Taylor expansion of the partial LOO-map ${\bm{f}}_{i}$, the resulting change in the $i$-th embedding point is expressed as
${\bm{f}}_{i}({\bm{x}}_{i}+\epsilon{\bm{e}})-{\bm{y}}_{i}=-\epsilon{\bm{H}}_{i}^{-1}\sum_{k:k\neq i}\frac{\partial^{2}\mathcal{L}(w({\bm{y}}_{i},{\bm{y}}_{k});v_{i,k}({\bm{X}}))}{\partial{\bm{y}}_{i}\partial{\bm{x}}_{k}^{\top}}{\bm{e}}+o(\epsilon),$ (8)
@@
where ${\bm{H}}_{i}$ denote the Hessian matrix of the partial LOO loss $L_{i}({\bm{y}};{\bm{x}}_{i})$ with respect to $\bm{y}$} at ${\bm{y}}={\bm{y}}_{i}$. Notably, when $\epsilon=0$ (no perturbation), we have ${\bm{f}}_{i}({\bm{x}}_{i})={\bm{y}}_{i}$. Denote the total loss as
$\mathfrak{L}({\bm{y}}_{1},\ldots,{\bm{y}}_{n};{\bm{X}})=\sum_{1\leq i<j\leq n}\mathcal{L}(w(\bm{y}_{i},\bm{y}_{j});v_{i,j}({\bm{X}}))+Z(\bm{Y}).$
Then, ${\bm{H}}_{i}$ can be written as
${\bm{H}}_{i}=\left.\frac{\partial^{2}L_{i}({\bm{y}};{\bm{x}}_{i})}{\partial{\bm{y}}\partial{\bm{y}}^{\top}}\right|_{{\bm{y}}={\bm{y}}_{i}}=\frac{\partial^{2}\mathfrak{L}({\bm{y}}_{1},\ldots,{\bm{y}}_{n};{\bm{X}})}{\partial{\bm{y}}_{i}\partial{\bm{y}}_{i}^{\top}},$
@@
i.e., ${\bm{H}}_{i}$ also equals to the Hessian matrix of the total loss $\mathfrak{L}$ with respect to the $i$-th variable taking value at ${\bm{y}}_{i}$.
@@
Importantly, ${\bm{H}}_{i}$ is independent of the perturbation direction $\bm{e}$}. The more singular ${\bm{H}}_{i}$ is, the more sensitive the embedding point of ${\bm{x}}_{i}$ becomes to infinitesimal perturbations. Thus, we define the singularity score of the $i$-th data point as the inverse of the smallest eigenvalue of the Hessian matrix of $\mathfrak{L}$, that is $\lambda_{\min}^{-1}({\bm{H}}_{i})$. Supplementary file Section E.1 provides detailed derivations of Eqn. 8 and Supplementary file Section E.2 provides expressions of singularity scores for t-SNE, UMAP and LargeVis.
A.3 Scoring Metrics and Statistical Tests
A.3.1 Entropy of class probabilities
@@
For a classification task, a statistical or machine learning algorithm outputs predicted class probabilities for a test data point. For example, in neural networks, the probabilities are typically obtained through a softmax operation in the final layer. Often, the model predicts a class with the largest probability among all classed. The entropy of the probabilities can quantify how confident the model is in its prediction.
@@
For a classification task of $k$ classes, if we denote the outputs class probabilities for one data point $\bm{x}$} as ${\bm{p}}=(p_{1},\ldots,p_{k})$, then we define the entropy as $E({\bm{p}})=-\sum_{j=1}^{k}p_{j}\log(p_{j})$. This quantity is widely used for measuring the class uncertainty.
A.3.2 Entropy difference
@@
We will describe an uncertainty measurement given access to the labels of input points. For a dataset $({\bm{x}}_{i})_{i\leq n}$ with clustering structures, we posit the following $k$-components Gaussian mixture model (GMM) which each ${\bm{x}}_{i}$ is sampled from. Consider a uniform prior on the $k$ clusters, i.e., $p(A_{j})=\frac{1}{k}$, $j=1,2,\ldots,k$. Given cluster membership $A_{j}$, we define the conditional probability density function
$p(\bm{x}|A_{j})=g(\bm{x}|\bm{\mu}_{j},\bm{\Sigma}_{j}),$
where $\bm{\mu}_{j}$, $\bm{\Sigma}_{j}$ are the mean and covariance matrix in the $j$-th component, and $g(\bm{x}|\bm{\mu}_{j},\bm{\Sigma}_{i})$, $j=1,2,\ldots,k$ are the Gaussian density functions with mean $\bm{\mu}_{j}$ and covariance matrix $\bm{\Sigma}_{j}$. We then have the posterior probability of $A_{j}$ given an observation $\bm{x}$ as
$p(A_{j}|\bm{x})=\frac{p(\bm{x}|A_{j})}{\sum_{j=1}^{k}p(\bm{x}|A_{j})}.$ (9)
In the analysis of neighbor embedding methods, we will use the posterior probabilities as an uncertainty measurement. Given the ground-truth labels of the data points, we can fit two GMMs, one in the input space and the other in the embedding space, yielding estimated parameters $({\bm{\mu}}_{j},\text{$\Sigma$}_{j})_{j\leq k}$ for each fitted GMM. Then we can calculate the posterior probabilities of each data point belonging to the $k$ components by Eqn. 9 with fitted parameters, in both the input space and the embedding space. For any data point, denote the posterior probabilities in input space as $\bm{p}=(p_{1},p_{2},\ldots,p_{k})$ and in embedding space as $\bm{q}=(q_{1},q_{2},\ldots,q_{k})$. Finally, we define the entropy difference for each point as the difference between the entropy of $\bm{p}$ and the entropy of $\bm{q}$, i.e., $E({\bm{p}})-E({\bm{q}})=-\sum_{j=1}^{k}p_{j}\log(p_{j})+\sum_{j=1}^{k}q_{j}\log(q_{j})$.
@@
The entropy difference measures the decreased uncertainty of cluster membership. A positive entropy difference means $E({\bm{q}})<E({\bm{p}})$, so the associated data point appears to be less ambiguous in cluster membership after embedding. Vice versa, a negative entropy difference means increased uncertainty of a data point.
@@
Since calculating entropy differences is based on the ground-truth labels and fitting a clear statistical model, we believe that entropy differences are a relatively objective evaluation of visual uncertainty. If a diagnosis score without label information is aligned with the entropy difference, then the diagnosis score is likely to be reliable.
A.3.3 Evaluation score of neighborhood preservation
@@
We calculate point-wise neighborhood preservation scores to evaluate how well the local structures are preserved by an embedding algorithm. Given the input matrix $\bm{X}$} and the embedding matrix $\bm{Y}$}, to calculate the neighborhood preservation score for the $i$-th point, we first identify its $k$-nearest neighbors in the input space, with their indices denoted as $\mathcal{N}_{i}=\{i_{1},i_{2},\ldots,i_{k}\}$. Then, we compute the distances from the $i$-th point to its neighbors in both the input and embedding spaces:
${\bm{d}}_{i}^{\text{input}}$ $=[d({\bm{x}}_{i},{\bm{x}}_{i_{1}}),\ldots,d({\bm{x}}_{i},{\bm{x}}_{i_{k}})]^{\top}$
${\bm{d}}_{i}^{\text{embedding}}$ $=[d({\bm{y}}_{i},{\bm{y}}_{i_{1}}),\ldots,d({\bm{y}}_{i},{\bm{y}}_{i_{k}})]^{\top}.$
@@
The neighborhood preservation score for the $i$-th point is defined as the correlation between ${\bm{d}}_{i}^{\text{input}}$ and ${\bm{d}}_{i}^{\text{embedding}}$. A higher correlation indicates better preservation of the neighborhood structure.
@@ A.3.4 Indices and metrics for evaluating clustering quality
Consider $m$ clusters and in each cluster $i$, there are $n_{i}$ data points, denoted as $\{{\bm{x}}_{ij}\}_{1\leq j\leq n_{i}}$. The centroid for each cluster is denoted as ${\bm{x}}_{i\cdot}=\frac{1}{n_{i}}\sum_{j=1}^{n_{i}}{\bm{x}}_{ij}$ and the mean of all data points is denoted as ${\bm{x}}_{\cdot\cdot}=\frac{1}{n}\sum_{i=1}^{m}\sum_{j=1}^{n_{i}}{\bm{x}}_{ij}$.
Denote the total sum of squares (TSS) and within-cluster sum of squares (WSS) by
$\text{TSS}=\sum_{i=1}^{m}\sum_{j=1}^{n_{i}}\|{\bm{x}}_{ij}-{\bm{x}}_{\cdot\cdot}\|_{2}^{2},\quad\text{WSS}=\sum_{i=1}^{m}\sum_{j=1}^{n_{i}}\|{\bm{x}}_{ij}-{\bm{x}}_{i\cdot}\|_{2}^{2}.$
@@
A.3.5 Statistical tests for distribution difference of singularity scores
We have claimed that embedding points with large singularity scores tend to appear in random locations under small perplexities, but appear in the periphery of clusters under large perplexity. To quantitatively verify such distinction, we conduct several statistical tests and find that the results of the tests support our claim about the distribution difference (see Supplementary Table S4). We provide the details of the tests.
@@ A.3.5 Statistical tests for distribution difference of singularity scores
Given the embedding ${\bm{Y}}=[{\bm{y}}_{1},\ldots,{\bm{y}}_{n}]^{\top}$ and the cluster label of each point as well as their singularity scores ${\bm{s}}=[s_{1},\ldots,s_{n}]^{\top}$, we can first calculate the distance of each point to its cluster center. The distance vector is denoted as ${\bm{d}}=[d_{1},\ldots,d_{n}]^{\top}$. We then conduct the Spearman’s rank correlation test [53] on the singularity scores $\bm{s}$} and the distances to cluster center $\bm{d}$}. The tests shows that there is no significant correlation under low perplexity but significant correlation under larger perplexity (see $p$-values in Supplementary Table S4).
@@
A.4 Data processing
A.4.1 Gaussian mixture data
A Gaussian mixture model with $k$ components is a linear combination of $k$ component Gaussian densities. The probability density function of the random variable $\bm{x}$ generated by Gaussian mixture model [55] is
$p(\bm{x}|\bm{\lambda})=\sum_{i=1}^{k}\pi_{i}g(\bm{x}|\bm{\mu}_{i},\bm{\Sigma}_{i}),$
where $\bm{\mu}_{i}$, $\bm{\Sigma}_{i}$ are the mean and covariance matrix in the $i$-th component, the scalars $\pi_{i}$, $i=1,2,\ldots,k$ are the mixture weights satisfying $\sum_{i=1}^{k}\pi_{i}=1$, and $g(\bm{x}|\bm{\mu}_{i},\bm{\Sigma}_{i})$, $i=1,2,\ldots,n$ are the probability density functions of the Gaussian distribution family with mean $\bm{\mu}_{i}$ and covariance matrix $\bm{\Sigma}_{i}$.
We generated Gaussian mixture data using the function rGMM in the R package MGMM (version 1.0.1.1).
A.4.2 Swiss roll data
@@
A.4.3 Deep learning feature data
@@
We used the pretrained ResNet-18 model [48] to perform a forward pass on the CIFAR-10 test dataset to extract features of dimension 512. We also performed the forward pass using the same pre-trained model on the Describable Textures Dataset (DTD) dataset [50] as our out-of-distribution data in Section 2.6. We also randomly subsampled both datasets to reduce computational load. Specifically, in Section 1, we sampled $5000$ images from the CIFAR-10 dataset as our deep learning feature data. In Section 2.6, we sampled $2000$ CIFAR-10 images and $1000$ DTD images, combining them into a dataset that includes OOD data points.
A.4.4 Mouse brain single-cell ATAC-seq data
@@
The ATAC-seq datasets have been preprocessed by Luecken et al. ([52]). We applied the R functions CreateSeuratObject, FindVariableFeatures and NormalizeData in R package Seurat to identify 1000 most variable genes for 3618 cells. The dataset was subsampled when being used to verify the LOO assumption.
A.4.5 Mouse embryonic stem cell differentiation data
@@
The single-cell RNA-seq data of mouse embryonic stem cells (mESCs) undergoing differentiation [51] were preprocessed, normalized, and scaled by following the standard procedures by R package Seurat using functions CreateSeuratObject, NormalizeData and ScaleData. We also used R function FindVariableFeatures to identify 2000 most variable genes for all 421 cells.
A.4.6 Human pancreatic tissue single-cell RNA-seq data
@@
The single-cell RNA-seq data generated from human pancreatic tissues [56] were preprocessed, normalized, and scaled by following the standard procedures described above. We also used R function FindVariableFeatures to identify 2000 most variable genes for all 2364 cells. The dataset was subsampled when being used to verify the LOO assumption.
@@
A.4.7 Single-cell RNA-seq data of PBMCs with treatment of interferon-beta
@@
The single-cell RNA-seq data generated from PBMCs treated with interferon-$\beta$ (IFNB) [57] were preprocessed, normalized, and scaled by following the standard procedures described above. We used R function FindVariableFeatures to identify 2000 most variable genes for all 6548 cells. The dataset was subsampled when being used to verify the LOO assumption.
A.5 Implementation of t-SNE
@@
We used the function Rtsne in the R package Rtsne (version 0.17) to perform the t-SNE algorithm. We choose theta = 0 to perform exact t-SNE. We also adjusted the code in Rtsne to access the similarity scores $(v_{i,j}({\bm{X}}))_{i<j}$. The adjusted function Rtsne can be found in https://github.com/zhexuandliu/NE-Reliability-MapContinuity.
Appendix B Data availability

### v2

@@ Methods
In general, perturbation scores are not sensitive to perturbation lengths. Supplementary Fig. S6 illustrates the perturbation scores of the CIFAR10 deep learning feature data for three perturbation lengths ($\lambda\in\{1,2,3\}$). Points with high perturbation scores remain consistent across different perturbation lengths. In practice, we recommend that users run perturbation scores on a subset of data points and test with a few different perturbation lengths. Conceptually, the perturbation score detects points that fall within a radius of $\lambda$ around the location of the OI discontinuity.
@@
Moreover, we provide two approximation algorithms to accelerate the calculation of the perturbation score for t-SNE along with a strategy for users to pre-screen points for which the perturbation score should be computed.
@@
For high-dimensional input data, often PCA as a pre-processing step is implemented before calculating the similarity scores. As similarity scores are recalculated for each perturbation we consider, PCA is repeated numerous times, leading to a significant increase in computation. Since PCA is robust to perturbing a single input point, we reuse the pre-processed input points after one PCA calculation based on the original input data. This approximation avoids multiple calculations of PCA. We find that this approximation is sufficiently accurate, as the differences between perturbation scores by approximation method 1 and the exact perturbation scores are empirically negligible (Supplementary Fig. S8a).
@@
Given the input data matrix ${\mathbf{X}}=[{\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{n}]^{\top}$ and perplexity $\mathcal{P}$, the computation of (exact) similarity scores $(v_{i,j}({\mathbf{X}}))_{i<j}$ in the t-SNE algorithm follows the steps below.
@@
    Calculate the pairwise distance $d_{ij}=\|{\mathbf{x}}_{i}-{\mathbf{x}}_{j}\|_{2}$ for $i,j=1,\ldots,n$.
@@
    $v_{i,j}({\mathbf{X}})=\frac{p_{j|i}+p_{i|j}}{2n}.$
The main computational bottleneck is at step 2, where we conduct a binary search algorithm for $n$ times to solve $(\sigma_{i})_{1\leq i\leq n}$.
To provide an approximation method, we note that when perturbing the $k$-th point, for $i\neq k$, Equation 7 still approximately holds for the original standard deviation $\sigma_{i}$ since only one of the terms has been changed. Therefore, we can set $\widetilde{\sigma}_{i}\approx\sigma_{i}$ for $i\neq k$ as an approximation to $(\widetilde{\sigma}_{i})_{1\leq i\leq n}$, the standard deviations after perturbation. In this way, we only need to conduct the binary search once to solve $\widetilde{\sigma}_{k}$, which significantly speeds up the calculation of the similarity scores after perturbation.
In terms of computational performance, approximation method 2 leads to a reduction of running time by nearly $80\%$ for a dataset of size $5000$. We also find that approximation method 2 is highly accurate. As shown in Supplementary Fig. S8b, perturbation scores based on approximation method 2 are approximately equal to the exact perturbation scores for most of the points.
@@
To further speed up the computation, we use the heuristic that embedding points receiving high perturbation scores are often found at the peripheries of clusters. This heuristic motivates us to calculate the perturbation scores only for the peripheral points in the embedding space, as these points are most likely to be unreliable. We find that applying this pre-screening step tends to find most of the unreliable points (Supplementary Fig. S7) with significantly increased computational speed.
@@
We use the function dbscan in the R package dbscan (version 1.2-0) to identify embeddings on the periphery of clusters.
Singularity score
@@
Given an input data matrix ${\mathbf{X}}=[{\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{n}]^{\top}$ and its embedding matrix ${\mathbf{Y}}=[{\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n}]^{\top}$, we describe our derivation of singularity scores. If we add an infinitesimal perturbation $\epsilon{\mathbf{e}}$ to ${\mathbf{x}}_{i}$, then by the Taylor expansion of the partial LOO-map ${\mathbf{f}}_{i}$, the resulting change in the $i$-th embedding point is expressed as
${\mathbf{f}}_{i}({\mathbf{x}}_{i}+\epsilon{\mathbf{e}})-{\mathbf{y}}_{i}=-\epsilon{\mathbf{H}}_{i}^{-1}\sum_{k:k\neq i}\frac{\partial^{2}\mathcal{L}(w({\mathbf{y}}_{i},{\mathbf{y}}_{k});v_{i,k}({\mathbf{X}}))}{\partial{\mathbf{y}}_{i}\partial{\mathbf{x}}_{k}^{\top}}{\mathbf{e}}+o(\epsilon),$ (8)
@@
where ${\mathbf{H}}_{i}$ denotes the Hessian matrix of the partial LOO loss $L_{i}({\mathbf{y}};{\mathbf{x}}_{i})$ with respect to $\mathbf{y}$} at ${\mathbf{y}}={\mathbf{y}}_{i}$. Notably, when $\epsilon=0$ (no perturbation), we have ${\mathbf{f}}_{i}({\mathbf{x}}_{i})={\mathbf{y}}_{i}$. Denote the total loss as
$\mathfrak{L}({\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n};{\mathbf{X}})=\sum_{1\leq i<j\leq n}\mathcal{L}(w({\mathbf{y}}_{i},{\mathbf{y}}_{j});v_{i,j}({\mathbf{X}}))+Z({\mathbf{Y}}).$
Then, ${\mathbf{H}}_{i}$ can be written as
${\mathbf{H}}_{i}=\left.\frac{\partial^{2}L_{i}({\mathbf{y}};{\mathbf{x}}_{i})}{\partial{\mathbf{y}}\partial{\mathbf{y}}^{\top}}\right|_{{\mathbf{y}}={\mathbf{y}}_{i}}=\frac{\partial^{2}\mathfrak{L}({\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n};{\mathbf{X}})}{\partial{\mathbf{y}}_{i}\partial{\mathbf{y}}_{i}^{\top}},$
@@
i.e., ${\mathbf{H}}_{i}$ is also equal to the Hessian matrix of the total loss $\mathfrak{L}$ with respect to the $i$-th variable taking value at ${\mathbf{y}}_{i}$.
@@
Importantly, ${\mathbf{H}}_{i}$ is independent of the perturbation direction $\mathbf{e}$}. The more singular ${\mathbf{H}}_{i}$ is, the more sensitive the embedding point of ${\mathbf{x}}_{i}$ becomes to infinitesimal perturbations. Thus, we define the singularity score of the $i$-th data point as the inverse of the smallest eigenvalue of the Hessian matrix of $\mathfrak{L}$, that is $\lambda_{\min}^{-1}({\mathbf{H}}_{i})$. Supplementary Methods A.1 provides detailed derivations of Equation 8 and Supplementary Methods A.2 provides expressions of singularity scores for t-SNE, UMAP and LargeVis.
Scoring Metrics and Statistical Tests
Entropy of class probabilities.
@@
For a classification task, a statistical or machine learning algorithm outputs predicted class probabilities for a test data point. For example, in neural networks, the probabilities are typically obtained through a softmax operation in the final layer. Often, the model predicts a class with the largest probability among all classes. The entropy of the probabilities can quantify how confident the model is in its prediction.
@@
For a classification task of $k$ classes, if we denote the outputs class probabilities for one data point $\mathbf{x}$} as ${\mathbf{p}}=(p_{1},\ldots,p_{k})$, then we define the entropy as $E({\mathbf{p}})=-\sum_{j=1}^{k}p_{j}\log(p_{j})$. This quantity is widely used for measuring class uncertainty.
Entropy difference.
@@
We will describe an uncertainty measurement given access to the labels of input points. For a dataset $({\mathbf{x}}_{i})_{i\leq n}$ with clustering structures, we posit the following $k$-component Gaussian mixture model (GMM) from which each ${\mathbf{x}}_{i}$ is sampled. Consider a uniform prior on the $k$ clusters, i.e., $p(A_{j})=\frac{1}{k}$, $j=1,2,\ldots,k$. Given cluster membership $A_{j}$, we define the conditional probability density function
[…]
$p(A_{j}|{\mathbf{x}})=\frac{p({\mathbf{x}}|A_{j})}{\sum_{j=1}^{k}p({\mathbf{x}}|A_{j})}.$ (9)
[…]
@@
The entropy difference measures the amount of decreased uncertainty of cluster membership. A positive entropy difference means $E({\mathbf{q}})<E({\mathbf{p}})$, so the associated data point appears to be less ambiguous in cluster membership after embedding. Vice versa, a negative entropy difference means increased uncertainty after embedding.
@@
Since calculating entropy differences is based on ground-truth labels and fitting a clear statistical model, we believe that entropy differences are a relatively objective evaluation of visual uncertainty. If a diagnostic score without label information is aligned with the entropy difference, then the diagnostic score is likely to be reliable.
Evaluation score of neighborhood preservation.
@@
We calculate point-wise neighborhood preservation scores to evaluate how well the local structures are preserved by an embedding algorithm. Given the input matrix $\mathbf{X}$} and the embedding matrix $\mathbf{Y}$}, to calculate the neighborhood preservation score for the $i$-th point, we first identify its $k$-nearest neighbors in the input space, with their indices denoted as $\mathcal{N}_{i}=\{i_{1},i_{2},\ldots,i_{k}\}$. Then, we compute the distances from the $i$-th point to its neighbors in both the input and embedding spaces:
${\mathbf{d}}_{i}^{\text{input}}$ $=[d({\mathbf{x}}_{i},{\mathbf{x}}_{i_{1}}),\ldots,d({\mathbf{x}}_{i},{\mathbf{x}}_{i_{k}})]^{\top}$
${\mathbf{d}}_{i}^{\text{embedding}}$ $=[d({\mathbf{y}}_{i},{\mathbf{y}}_{i_{1}}),\ldots,d({\mathbf{y}}_{i},{\mathbf{y}}_{i_{k}})]^{\top}.$
@@
The neighborhood preservation score for the $i$-th point is defined as the correlation between ${\mathbf{d}}_{i}^{\text{input}}$ and ${\mathbf{d}}_{i}^{\text{embedding}}$. A higher correlation indicates better preservation of the neighborhood structure.
@@
Consider $m$ clusters and in each cluster $i$, there are $n_{i}$ data points, denoted as $\{{\mathbf{x}}_{ij}\}_{1\leq j\leq n_{i}}$. The centroid for each cluster is denoted as ${\mathbf{x}}_{i\cdot}=\frac{1}{n_{i}}\sum_{j=1}^{n_{i}}{\mathbf{x}}_{ij}$ and the mean of all data points is denoted as ${\mathbf{x}}_{\cdot\cdot}=\frac{1}{n}\sum_{i=1}^{m}\sum_{j=1}^{n_{i}}{\mathbf{x}}_{ij}$.
Denote the total sum of squares (TSS) and the within-cluster sum of squares (WSS) by
$\text{TSS}=\sum_{i=1}^{m}\sum_{j=1}^{n_{i}}\|{\mathbf{x}}_{ij}-{\mathbf{x}}_{\cdot\cdot}\|_{2}^{2},\quad\text{WSS}=\sum_{i=1}^{m}\sum_{j=1}^{n_{i}}\|{\mathbf{x}}_{ij}-{\mathbf{x}}_{i\cdot}\|_{2}^{2}.$
@@
Statistical tests for distribution difference of singularity scores.
We have claimed that embedding points with large singularity scores tend to appear in random locations under small perplexities but appear in the periphery of clusters under large perplexities. To quantitatively verify such distinction, we conduct several statistical tests and find that the results of the tests support our claim about the distribution difference (see Supplementary Table S5). We provide the details of the tests as follows.
@@
Given the embedding ${\mathbf{Y}}=[{\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n}]^{\top}$ and the cluster label of each point as well as their singularity scores ${\mathbf{s}}=[s_{1},\ldots,s_{n}]^{\top}$, we can first calculate the distance of each point to its cluster center. The distance vector is denoted as ${\mathbf{d}}=[d_{1},\ldots,d_{n}]^{\top}$. We then conduct the Spearman’s rank correlation test [56] on the singularity scores $\mathbf{s}$} and the distances to cluster center $\mathbf{d}$}. The tests show that there is no significant correlation under low perplexity but a significant correlation under larger perplexity (see $p$-values in Supplementary Table S5).
@@
Benchmark Methods for OOD Detection
Kernel PCA.
We implemented the state-of-the-art kernel PCA method for out-of-distribution detection [52] to benchmark against the perturbation score. Since our perturbation score does not require separate training and testing steps and was directly applied to the dataset, kernel PCA was trained on the dataset and then evaluated on the same dataset to ensure a fair comparison. Additionally, to maintain consistency with the default PCA preprocessing step in the t-SNE algorithm, we applied PCA before training, retaining the first 50 principal components.
One-class SVM.
We implemented the one-class SVM [53] using the OneClassSVM function from the Python package scikit-learn, employing a polynomial kernel for optimal performance. Since our perturbation score does not require separate training and testing steps and was directly applied to the dataset, one-class SVM was also trained on the dataset and then evaluated on the same dataset to ensure a fair comparison. To align with the preprocessing step in t-SNE, we first applied PCA, reducing the data to its top 50 principal components before training.
Datasets
Gaussian mixture data.
A Gaussian mixture model with $k$ components is a linear combination of $k$-component Gaussian densities. The probability density function of the random variable $\mathbf{x}$} generated by Gaussian mixture model [58] is
[…]
We randomly generated Gaussian mixture datasets with various numbers of components and mixture weights using the function rGMM in the R package MGMM (version 1.0.1.1).
Swiss roll data.
@@
We randomly generated Swiss roll datasets and used the function Rtsne in the R package Rtsne (version 0.17) to obtain the t-SNE embeddings of the datasets. We computed the perturbation scores with perturbation length $1$ in Fig. 7b.
Deep learning feature data.
@@
We used the pretrained ResNet-18 model to perform a forward pass on the CIFAR-10 dataset to extract features of dimension 512. We also performed the forward pass using the same pre-trained model on the Describable Textures Dataset (DTD) dataset [51] as our out-of-distribution data in Fig. 5. We also randomly subsampled both datasets to reduce computational load. Specifically, in Fig. 1, we sampled $5000$ images from the CIFAR-10 test dataset as our deep learning feature data and obtained the t-SNE embedding under perplexity $125$. We then computed the perturbation scores with perturbation length $2$. In Fig. 5, we sampled $2000$ CIFAR-10 images and $1000$ DTD images, combining them into a dataset that includes OOD data points. We obtained the t-SNE embedding under perplexity $100$ and computed the perturbation scores with perturbation length $2$.
Mouse brain single-cell ATAC-seq data.
@@
The ATAC-seq dataset was created to capture the gene activity of mouse brain cells. The dataset has been preprocessed by Luecken et al. [59]. We applied the R functions CreateSeuratObject, FindVariableFeatures and NormalizeData in R package Seurat to identify 1000 most variable genes for 3618 cells. The dataset was subsampled when being used to verify the LOO assumption.
Mouse embryonic stem cell differentiation data.
@@
The single-cell RNA-seq dataset was constructed to investigate the dynamics of gene expression of mouse embryonic stem cells (mESCs) undergoing differentiation [54]. The dataset was preprocessed, normalized, and scaled by following the standard procedures by R package Seurat using functions CreateSeuratObject, NormalizeData and ScaleData. We also used R function FindVariableFeatures to identify the 2000 most variable genes for all 421 cells.
Human pancreatic tissue single-cell RNA-seq data.
@@
The single-cell RNA-seq data generated from human pancreatic tissues [60] provides a comprehensive view of gene expression across 8 different cell types in pancreatic tissue. The dataset was preprocessed, normalized, and scaled by following the standard procedures described above. We also used R function FindVariableFeatures to identify the 2000 most variable genes for all 2364 cells. The dataset was subsampled when being used to verify the LOO assumption.
@@
Single-cell RNA-seq data of PBMCs with treatment of interferon-beta.
@@
This single-cell RNA-seq dataset profiles gene expression in peripheral blood mononuclear cells (PBMCs) following interferon-$\beta$ (IFNB) treatment, capturing cellular responses to immune stimulation [61]. The dataset was preprocessed, normalized, and scaled by following the standard procedures described above. We used R function FindVariableFeatures to identify the 2000 most variable genes for all 6,548 cells. The dataset was subsampled when being used to verify the LOO assumption.
Mouse mammary epithelial single-cell data.
This dataset contains the gene expression profile of mammary epithelial cells across from two mice at four developmental stages: nulliparous, mid-gestation, lactation, and post-involution [55]. The dataset was preprocessed, normalized, and scaled by following the standard procedures described above. We used R function FindVariableFeatures to identify the 2000 most variable genes for all 25,806 cells.
Implementation of t-SNE
@@
We used the function Rtsne in the R package Rtsne (version 0.17) to perform the t-SNE algorithm. We choose theta = 0 to perform exact t-SNE. We also adjusted the code in Rtsne to access the similarity scores $(v_{i,j}({\mathbf{X}}))_{i<j}$. The adjusted function Rtsne can be found in https://github.com/zhexuandliu/MapContinuity-NE-Reliability.
Data availability